The total cholesterol levels of a sample of men aged​ 35-44 are normally distributed with a mean of 208 milligrams per deciliter and a standard deviation of 36.6 milligrams per deciliter. ​(a) What percent of the men have a total cholesterol level less than 214 milligrams per deciliter of​ blood? ​(b) If 246 men in the​ 35-44 age group are randomly​ selected, about how many would you expect to have a total cholesterol level greater than 265 milligrams per deciliter of​ blood? ) The percent of the men that have a total cholesterol level less than 214 milligrams per deciliter of blood is %. ​(Round to two decimal places as​ needed.) ​(b) Of the 246 men​ selected, nothing would be expected to have a total cholesterol level greater than 265 milligrams per deciliter of blood. ​(Round to the nearest integer as​ needed.)
Answer #1

Let X be the cholestrol level of men

X~N(208,36.6)

a)

=0.5+0.0636=0.5636

=56.36%

b)

=0.5-0.4406 =0.0594

If 246 men, the number of men who have cholestrol level greater than 265 is = 246*0.0594 =14.6115

Answer: 15